# A Survey on 3D Gaussian Splatting

### v7

@@
Abstract
3D Gaussian splatting (GS) has emerged as a transformative technique in explicit radiance field and computer graphics. This innovative approach, characterized by the use of millions of learnable 3D Gaussians, represents a significant departure from mainstream neural radiance field approaches, which predominantly use implicit, coordinate-based models to map spatial coordinates to pixel values. 3D GS, with its explicit scene representation and differentiable rendering algorithm, not only promises real-time rendering capability but also introduces unprecedented levels of editability. This positions 3D GS as a potential game-changer for the next generation of 3D reconstruction and representation. In the present paper, we provide the first systematic overview of the recent developments and critical contributions in the domain of 3D GS. We begin with a detailed exploration of the underlying principles and the driving forces behind the emergence of 3D GS, laying the groundwork for understanding its significance. A focal point of our discussion is the practical applicability of 3D GS. By enabling unprecedented rendering speed, 3D GS opens up a plethora of applications, ranging from virtual reality to interactive media and beyond. This is complemented by a comparative analysis of leading 3D GS models, evaluated across various benchmark tasks to highlight their performance and practical utility. The survey concludes by identifying current challenges and suggesting potential avenues for future research. Through this survey, we aim to provide a valuable resource for both newcomers and seasoned researchers, fostering further exploration and advancement in explicit radiance field.
Index Terms:
3D Gaussian Splatting, Explicit Radiance Field, Real-time Rendering, Scene Understanding
1 Introduction
@@
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods [1, 2, 3]. These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion [4] and subsequent advancements in multi-view stereo [5] algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and texture loss. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented fidelity. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive [6, 7, 8, 9], often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
Figure 1: The number of published papers and official GitHub stars on 3D GS. The set of statistics is sourced from # Papers and # GitHub Stars.
@@
It is in this context that 3D Gaussian splatting (GS) [10] emerges, not merely as an incremental improvement but as a paradigm-shifting approach that redefines the boundaries of scene representation and rendering. While NeRF excelled in creating photorealistic images, the need for faster, more efficient rendering methods was becoming increasingly apparent, especially for applications (e.g., virtual reality and autonomous driving) that are highly sensitive to latency. 3D GS addressed this need by introducing an advanced, explicit scene representation that models a scene using millions of learnable 3D Gaussians in space. Unlike the implicit, coordinate-based models [11, 12], 3D GS employs an explicit representation and highly parallelized workflows, facilitating more efficient computation and rendering. The innovation of 3D GS lies in its unique blend of the benefits of differentiable pipelines and point-based rendering techniques [13, 14, 15, 16, 17]. By representing scenes with learnable 3D Gaussians, it preserves the strong fitting capability of continuous volumetric radiance fields, essential for high-quality image synthesis, while simultaneously avoiding the computational overhead associated with NeRF based methods (e.g., computationally expensive ray-marching, and unnecessary calculations in empty space).
@@
The introduction of 3D GS is not just a technical advancement; it represents a fundamental shift in how we approach scene representation and rendering in computer vision and graphics. By enabling real-time rendering capabilities without compromising on visual quality, 3D GS opens up a plethora of possibilities for applications ranging from virtual reality and augmented reality to real-time cinematic rendering and beyond [18, 19, 20, 21]. This technology holds the promise of not only enhancing existing applications but also enabling new ones that were previously unfeasible due to computational constraints. Furthermore, 3D GS’s explicit scene representation offers unprecedented flexibility to control the objects and scene dynamics, a crucial factor in complex scenarios involving intricate geometries and varying lighting conditions [22, 23, 24]. This level of editability, combined with the efficiency of the training and rendering process, positions 3D GS as a transformative force in shaping future developments in relevant fields.
§2.1§2.2§3.1§3.2§4.1§4.2§4.3§4.4§4.5§4.6§4.7§5.1§5.2§5.3§5.4§5.5§5.6§5.7§6.1§6.2§6.3§6.4§6.5§1§2§3§4§5§6§7§8
Figure 2: Structure of the overall review.
@@
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that 3D GS is a very recent innovation (Fig. 1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. The selected follow-up works are primarily sourced from top-tier conferences, to provide a thorough and up-to-date (Dec. 2024) analysis of the theoretical foundations, remarkable developments, and burgeoning applications of 3D GS. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain. The distinctions of this survey from existing literature [25, 26, 27, 28] are evident in the following aspects:
@@
2 Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), including both implicit and explicit ones. Sec. 2.2 further establishes linkages with relevant rendering algorithms and terminologies. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys [29, 30, 31, 32, 33] for more insights.
2.1 Radiance Field
$\bullet$ Implicit Radiance Field. An implicit radiance field represents light distribution in a scene without explicitly defining the geometry of the scene. In the deep learning era, neural networks are often used to learn a continuous volumetric scene representation [34, 35]. The most prominent example is NeRF [12]. In NeRF (Fig. 3a), one or more MLPs are used to map a set of spatial coordinates $(x,y,z)$ and viewing directions $(\theta,\phi)$ to color $c$ and volume density $\sigma$:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{MLP}(x,y,z,\theta,\phi)$}. (1)
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load due to volumetric ray marching. Note that typically, the color $c$ is direction-dependent, whereas the volume density $\sigma$ is not [12].
$\bullet$ Explicit Radiance Field. An explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points [36, 37]. Each element in this structure stores the radiance information for its respective location. This allows for direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. Similar to the implicit radiance field, the explicit one is written as:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{DataStructure}(x,y,z,\theta,\phi)$}, (2)
@@ 2.1 Radiance Field
$\bullet$ 3D Gaussian Splatting: Best-of-Both Worlds. 3D GS [10] is an explicit radiance field with the advantages of implicit radiance fields. Concretely, it leverages the strengths of both paradigms by utilizing learnable 3D Gaussians as the basis elements of DataStructure. Note that 3D GS encodes the opacity $\alpha$ directly for each Gaussian, as opposed to approaches of first establishing density $\sigma$ and then computing opacity based on that density. As in previous reconstruction work, 3D Gaussians are optimized under the supervision of multi-view images to represent the scene. Such a 3D Gaussian based differentiable pipeline combines the benefits of neural network based optimization and explicit, structured data storage. This hybrid approach aims to achieve real-time, high-quality rendering and requires less training time, particularly for complex scenes and high-resolution outputs.
2.2 Context and Terminology
@@
$\bullet$ Volumetric rendering aims to transform a 3D volumetric representation into an image by integrating radiance along camera rays. A camera ray $\bm{r}(t)$ can be parameterized as: $\bm{r}(t)\!=\!\bm{o}\!+\!t\bm{d},t\!\in\![t_{\text{near}},t_{\text{far}}],$ where $\bm{o}$ represents the ray origin (camera center), $\bm{d}$ is the ray direction, and $t$ indicates the distance along the ray between near and far clipping planes. The pixel color $C(\bm{r})$ is computed through a line integral along the ray $\bm{r}(t)$, mathematically expressed as [12]:
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C(\bm{r})=\int_{t_{\text{near}}}^{t_{\text{far}}}T(t)\,\sigma(\bm{r}(t))\,c(\bm{r}(t),\bm{d})\,dt,$} (3)
@@
where $\sigma(\bm{r}(t))$ is the volume density at point $\bm{r}(t)$, $c(\bm{r}(t),\bm{d})$ is the color at that point, and $T(t)$ is the transmittance. Ray-marching directly approximates the volumetric rendering integral by systematically “stepping” along a ray and sampling the scene’s properties at discrete intervals. NeRF [12] shares the same spirit of ray-marching and introduces importance sampling and positional encoding to improve the quality of synthesized images. While providing high-quality results, ray-marching is computationally expensive, especially for high-resolution images.
@@
$\bullet$ Point-based rendering represents another class of rendering algorithms, of which 3D GS introduces a notable implementation. Its simplest form [38] rasterizes point clouds with a fixed size, which introduces drawbacks such as holes and rendering artifacts. Seminal works addressed these limitations through various methods, including: i) splatting point primitives with a spatial extent [14, 39, 40, 15], and ii) more recently, embedding neural features directly into points for subsequent network-based rendering [41, 42]. 3D GS uses 3D Gaussian as the point primitive that contains explicit attributes (e.g., color and opacity) instead of implicit neural features. The rendering approach, i.e., point-based $\alpha$-blending (exemplified in Eq. 5), shares the same image formation model as NeRF-style volumetric rendering (Eq. 3[10], but demonstrates substantial speed advantages. This advantage originates from fundamental algorithmic differences. NeRFs approximate a line integral along a ray for each pixel, requiring expensive sampling. Point-based methods render point clouds using rasterization, which inherently benefits from parallel computational strategies [43].
3 3D Gaussian Splatting: Principles
@@
3.1 Rendering with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching, sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources [10]. By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” [39, 40] (see Fig. 3b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 3, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.

### v8

@@
Abstract.
3D Gaussian splatting (GS) has emerged as a transformative technique in radiance fields. Unlike mainstream implicit neural models, 3D GS uses millions of learnable 3D Gaussians for an explicit scene representation. Paired with a differentiable rendering algorithm, this approach achieves real-time rendering and unprecedented editability, making it a potential game-changer for 3D reconstruction and representation. In the present paper, we provide the first systematic overview of the recent developments and critical contributions in 3D GS. We begin with a detailed exploration of the underlying principles and the driving forces behind the emergence of 3D GS, laying the groundwork for understanding its significance. A focal point of our discussion is the practical applicability of 3D GS. By enabling unprecedented rendering speed, 3D GS opens up a plethora of applications, ranging from virtual reality to interactive media and beyond. This is complemented by a comparative analysis of leading 3D GS models, evaluated across various benchmark tasks to highlight their performance and practical utility. The survey concludes by identifying current challenges and suggesting potential avenues for future research. Through this survey, we aim to provide a valuable resource for both newcomers and seasoned researchers, fostering further exploration and advancement in explicit radiance field.
3D Gaussian Splatting, Explicit Radiance Field
copyright: acmcopyrightccs: General and reference Surveys and overviewsccs: Computing methodologies Renderingccs: Computing methodologies Computer vision
1. Introduction
@@
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods (Gortler et al., 2023; Levoy and Hanrahan, 2023; Buehler et al., 2023). These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion (Snavely et al., 2006) and subsequent advancements in multi-view stereo (Goesele et al., 2007) algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and texture loss. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented fidelity. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive (Garbin et al., 2021; Müller et al., 2022), often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
@@
It is in this context that 3D Gaussian splatting (GS) (Kerbl et al., 2023) emerges, not merely as an incremental improvement but as a paradigm-shifting approach that redefines the boundaries of scene representation and rendering. While NeRF excelled in creating photorealistic images, the need for faster, more efficient rendering methods was becoming increasingly apparent, especially for applications (e.g., virtual reality and autonomous driving) that are highly sensitive to latency. 3D GS addressed this need by introducing an advanced, explicit scene representation that models a scene using millions of learnable 3D Gaussians in space. Unlike the implicit, coordinate-based models (Sitzmann et al., 2019; Mildenhall et al., 2020), 3D GS employs an explicit representation and highly parallelized workflows, facilitating more efficient computation and rendering. The innovation of 3D GS lies in its unique blend of the benefits of differentiable pipelines and point-based rendering techniques (Pfister et al., 2000; Zwicker et al., 2001b; Ren et al., 2002; Yifan et al., 2019). By representing scenes with learnable 3D Gaussians, it preserves the strong fitting capability of continuous volumetric radiance fields, essential for high-quality image synthesis, while simultaneously avoiding the computational overhead associated with NeRF based methods (e.g., computationally expensive ray-marching, and unnecessary calculations in empty space).
The introduction of 3D GS is not just a technical advancement; it represents a fundamental shift in how we approach scene representation and rendering in computer vision and graphics. By enabling real-time rendering capabilities without compromising on visual quality, 3D GS opens up a plethora of possibilities for applications ranging from virtual reality
Figure 1. The number of published papers and official GitHub stars on 3D GS. The set of statistics is sourced from # Papers and # GitHub Stars.
@@
and augmented reality to real-time cinematic rendering and beyond (Kalkofen et al., 2008; Patney et al., 2016; Albert et al., 2017; Jiang et al., 2024c). This technology holds the promise of not only enhancing existing applications but also enabling new ones that were previously unfeasible due to computational constraints. Furthermore, 3D GS’s explicit scene representation offers unprecedented flexibility to control the objects and scene dynamics, a crucial factor in complex scenarios involving intricate geometries and varying lighting conditions (Lu et al., 2024b; Saito et al., 2024; Zhang et al., 2024e). This level of editability, combined with the efficiency of the training and rendering process, positions 3D GS as a transformative force in shaping future developments in relevant fields.
@@
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that 3D GS is a very recent innovation (c.f. Fig. 1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. We systematically review the foundational period of 3D GS, with selected works primarily sourced from top-tier conferences. Our analysis concentrates on the initial explosion of research following its introduction in 2023 through 2024, covering the theoretical underpinnings, landmark developments, and early applications that shaped the field. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain. The distinctions of this survey from existing literature (Fei et al., 2024; Dalal et al., 2024; Bao et al., 2024; Wu et al., 2024e) are evident in the following aspects:
@@
2. Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), including both implicit and explicit ones. Sec. 2.2 further establishes linkages with relevant rendering algorithms and terminologies. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys (Kobbelt and Botsch, 2004; Xie et al., 2022; Wang et al., 2024b; Tewari et al., 2022; Han et al., 2019) for more insights.
§1§2§3§4§5§6§7§8§2.1§2.2§3.1§3.2§4.1§4.2§4.3§4.4§4.5§4.6§4.7§5.1§5.2§5.3§5.4§5.5§5.6§5.7§6.1§6.2§6.3§6.4§6.5
Figure 2. Structure of the overall review.
2.1. Radiance Field
$\bullet$ Implicit Radiance Field. An implicit radiance field represents light distribution in a scene without explicitly defining the geometry of the scene. In the deep learning era, neural networks are often used to learn a continuous volumetric scene representation (Mescheder et al., 2019; Park et al., 2019). The most prominent example is NeRF (Mildenhall et al., 2020). In NeRF (Fig. 4a), one or more MLPs are used to map a set of spatial coordinates $(x,y,z)$ and viewing directions $(\theta,\phi)$ to color $c$ and volume density $\sigma$:
(1) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{MLP}(x,y,z,\theta,\phi)$}.
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load due to volumetric ray marching. Note that typically, the color $c$ is direction-dependent, whereas the volume density $\sigma$ is not (Mildenhall et al., 2020).
$\bullet$ Explicit Radiance Field. An explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points (Sun et al., 2022; Fridovich-Keil et al., 2023). Each element in this structure stores the radiance information for its respective location. This allows for direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. Similar to the implicit radiance field, the explicit one is written as:
(2) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{DataStructure}(x,y,z,\theta,\phi)$},
@@ 2.1. Radiance Field
$\bullet$ 3D Gaussian Splatting: Best-of-Both Worlds. 3D GS (Kerbl et al., 2023) is an explicit radiance field with the advantages of implicit radiance fields. Concretely, it leverages the strengths of both paradigms by utilizing learnable 3D Gaussians as the basis elements of DataStructure. Note that 3D GS encodes the opacity $\alpha$ directly for each Gaussian, as opposed to approaches of first establishing density $\sigma$ and then computing opacity based on that density. As in previous reconstruction work, 3D Gaussians are optimized under the supervision of multi-view images to represent the scene. Such a 3D Gaussian based differentiable pipeline combines the benefits of neural network based optimization and explicit, structured data storage. This hybrid approach aims to achieve real-time, high-quality rendering and requires less training time, particularly for complex scenes and high-resolution outputs.
2.2. Context and Terminology
@@
$\bullet$ Volumetric rendering aims to transform a 3D volumetric representation into an image by integrating radiance along camera rays. A camera ray $\bm{r}(t)$ can be parameterized as: $\bm{r}(t)\!=\!\bm{o}\!+\!t\bm{d},t\!\in\![t_{\text{near}},t_{\text{far}}],$ where $\bm{o}$ represents the ray origin (camera center), $\bm{d}$ is the ray direction, and $t$ indicates the distance along the ray between near and far clipping planes. The pixel color $C(\bm{r})$ is computed through a line integral along the ray $\bm{r}(t)$, mathematically expressed as (Mildenhall et al., 2020):
@@
(3) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C(\bm{r})=\int_{t_{\text{near}}}^{t_{\text{far}}}T(t)\,\sigma(\bm{r}(t))\,c(\bm{r}(t),\bm{d})\,dt,$}
@@
where $\sigma(\bm{r}(t))$ is the volume density at point $\bm{r}(t)$, $c(\bm{r}(t),\bm{d})$ is the color at that point, and $T(t)$ is the transmittance. Ray-marching directly approximates the volumetric rendering integral by systematically “stepping” along a ray and sampling the scene’s properties at discrete intervals. NeRF (Mildenhall et al., 2020) shares the same spirit of ray-marching and introduces importance sampling and positional encoding to improve the quality of synthesized images. While providing high-quality results, ray-marching is computationally expensive, especially for high-resolution images.
@@
$\bullet$ Point-based rendering represents another class of rendering algorithms, of which 3D GS introduces a notable implementation. Its simplest form (Grossman and Dally, 1998) rasterizes point clouds with a fixed size, which introduces drawbacks such as holes and rendering artifacts. Seminal works addressed these limitations through various methods, including: i) splatting point primitives with a spatial extent (Zwicker et al., 2001b, a, 2002; Ren et al., 2002), and ii) more recently, embedding neural features directly into points for subsequent network-based rendering (Aliev et al., 2020; Rückert et al., 2022). 3D GS uses 3D Gaussian as the point primitive that contains explicit attributes (e.g., color and opacity) instead of implicit neural features. The rendering approach, i.e., point-based $\alpha$-blending (exemplified in Eq. 5), shares the same image formation model as NeRF-style volumetric rendering (Eq. 3) (Kerbl et al., 2023), but demonstrates substantial speed advantages. This advantage originates from fundamental algorithmic differences. NeRFs approximate a line integral along a ray for each pixel, requiring expensive sampling. Point-based methods render point clouds using rasterization, which inherently benefits from parallel computational strategies (Lassner and Zollhofer, 2021).
3. 3D Gaussian Splatting: Principles
$C_{1}\penalty 10000\ \penalty 10000\ C_{2}$$C_{3}\penalty 10000\ \penalty 10000\ C_{4}$$C_{1}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{1}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{1}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{1}^{\prime}})$$C_{2}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{2}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{2}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{2}^{\prime}})$$C_{3}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{3}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{3}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{3}^{\prime}})$$C_{4}\!=\!{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{4}^{\prime}}c_{1}+{\color[rgb]{0.36328125,0.41796875,0.28125}\definecolor[named]{pgfstrokecolor}{rgb}{0.36328125,0.41796875,0.28125}\alpha_{4}^{\prime}}c_{2}(1-{\color[rgb]{0.6953125,0.32421875,0.2890625}\definecolor[named]{pgfstrokecolor}{rgb}{0.6953125,0.32421875,0.2890625}\alpha_{4}^{\prime}})$
Figure 3. An illustration of the forward process of 3D GS (see Sec. 3.1). (a) The splatting step projects 3D Gaussians into image space. (b) 3D GS divides the image into multiple non-overlapping patches, i.e., tiles. (c) 3D GS replicates the Gaussians which cover several tiles, assigning each copy an identifier, i.e., a tile ID. (d) By rendering the sorted Gaussians, we can obtain all pixels within the tile. Note that the computational workflows for pixels and tiles are independent and can be done in parallel. Best viewed in color.
@@
3.1. Rendering with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching,
$x,y,z,\theta,\phi$
$c\penalty 10000\ \&\penalty 10000\ \sigma$
Figure 4. NeRFs vs. 3D GS. (a) NeRF samples along the ray and then queries the MLP to obtain colors and densities, which can be seen as a backward mapping (ray tracing). (b) 3D GS projects all 3D Gaussians into the image space (i.e., splatting) and then performs parallel rendering, which can be viewed as a forward mapping (rasterization).
sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources (Kerbl et al., 2023). By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” (Zwicker et al., 2001a, 2002) (c.f. Fig. 4b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 4, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.